OpenStudy (anonymous):

Write the slope intercept form of the equation for the line for (-4), (-2); 4, 3. Please help....

OpenStudy (anonymous):

you can do this by finding the slope, plugging into point-slope form, and then conveting to slope intercept form

OpenStudy (anonymous):

do you need help?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

ok i'll give you the steps and you plug in to do the math. so first we're going to do the slope formula; (y2-y1)/(x2-x1)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

let me know when you have your answer

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

I think you need to double check your math.

OpenStudy (anonymous):

3 -(-2)/4-(-4) = 1/8

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sorry/ this answer should be 5/8

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

ok now we're going to take one point, and the slope and substitute it to point-slope formula do you know how to do this?

OpenStudy (anonymous):

no all of this is new to me

OpenStudy (anonymous):

oh dear ok, well no problem you're off to a great start point-slope formula is : y-y1=m(x-x1) x1,y1 = a coordinate m = slope

OpenStudy (anonymous):

can you figure out where substitution would come in?

OpenStudy (anonymous):

sorry, I don't really know how to do this

OpenStudy (anonymous):

it's ok. so we for the sake of simplicity we're going to use (4,3) for our coordinate, and 5/8 for our slope

OpenStudy (anonymous):

can you show me what the equation would look like when we substitute (4,3) to x1,y1 and 5/8 to me? y-y1=m(x-x1) the signs in the equations stay minus unless it is a negative number then it would have to change into addition sign bc (-)( -) = +. are you following me?

OpenStudy (anonymous):

I'm sorry, No I can't. I'm really lost

OpenStudy (anonymous):

it's ok i'm a visual learner too. y-3=5/8(x-4)

OpenStudy (anonymous):

now we can isolate this equation to slope intercept form

OpenStudy (anonymous):

thank you. Ok

OpenStudy (anonymous):

ok it's a little long because fractions but i will try my best to show you and explain

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

y-3=5/8(x-4) Distribute the (x-4) y-3=5/8x - 20/8 Add three to both sides y = 5/8x - 20/8 + 3 In order to add the like terms (-20/8) and 3 we must make a common denominator -20/8 can be simplified to -5/2. If we multiply 3 with 2/2 we will get 6/2. Now that we have a common denominator we will combine -5/2 and 6/2. This will give us 1/2. SO our final equation should be: y=5/8x+1/2

OpenStudy (anonymous):

I'm not sure if you've learned point-slope for yet but if not then don't sweat it! I'm in alg 2 so I probably used a different concept than what your current math class is teaching. Hopefully the visual helped you see how I arrived to my answer even if it's a little different than what you're used to :)

OpenStudy (anonymous):

I am trying to follow what you have wrote. Thank you

OpenStudy (anonymous):

you're welcome is that all you need help with?

OpenStudy (anonymous):

I meant to say no thank you. I will use your example for the rest of my problems. Have a good day

OpenStudy (anonymous):

You have helped me a lot. Thank you very much.

OpenStudy (anonymous):

you're welcome fren! good luck with your math course!